I have to do some problems like these two. Can anyone help me threw the steps on both of these so that I know how to do them.
example 1
Solve the system.
x + y = 9
x – y = 3

Example 2
Solve the system by addition.
–4x + 5y = 1
x – 4y = 19







Answer by wuwei96815(245)

These types of problems are best solve using elimination and substitution.

In the elimination part we have to eliminate one of the varibles so that we can solve for the other. When we add x+y=9 and x-y=3, we get 2x =12. What happened to the y's? They were eliminated (y-y=0). Now you can easily solve for x.
2x = 12
x = 6

Since we have the value of one of the variables, we can use substituion to find the other.
x+y=9
(6)+ y = 9
y=9+6 = 15

The second problem is more difficult because simply adding the equations together will not eliminate either variable. In the second problem try multiplying the second equation by 4 and then proceeding. Let me know how you make out.